Sunday, 30 May 2010

Having a drink

Queen Check

Queen cells still uncapped today and lava don't look too big, I think it's 8 days after capping!

Nuc with John's QC, both cells are empty, just had a quick peek and left.

Friday, 28 May 2010

Queen rearing`

Made original hive queenless 4 days ago, gave frame with young lave a day later. Today reunited colony, 4/5 QC which appear small but are uncapped. need to check in a few days

Sunday, 23 May 2010

New nuc

Placed johns QC in on Friday night, was going to leave them sealed for 5 days somewhere cool but yesterday the were buzzing so much I moved them to bottom of garden and opened. Weather has been very hot, late twenties

Thursday, 20 May 2010

First Q cells

John phoned today, 7 QC, and he offered me one. Made a small nuc tonight from the middle one,
one frame with sealed brood with a frame with stores on either side. The warm way, shook a few more bees in, sealed them up and placed somewhere cool ( the weather at the moment is very hot, 20+) put a feeder on.

Sunday, 16 May 2010

TReating varroa

Turned over the Apigaurd to expose more thymol to the brood yesterday, big drop today. MOved last brood above excluder yesterday ready for Queen rearing.

Tuesday, 11 May 2010

Cold weather

Even though weather is poor and day temp has not reached 10 degrees today the Apiguard is causing big mite fall. Need to get ready for queen production and nucs.

Tuesday, 4 May 2010

Another check

Right colony is a lot bigger than left, both a lot slower build up compared to last year.
There were a few varroa on monitoring boards so will now check over a week. New comb seems very yellow!
I am going to turn colonies so all will get morning sun and I can still access monitoring boards.